Logiweb bibliography of logic

Up Help

[0] logic

[1] base

[2] peano

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-07-04.UTC:07:55:10.732497 = MJD-53555.TAI:07:55:42.732497 = LGT-4627180542732497e-6